# Deep Fourier-embedded Network for RGB and Thermal Salient Object Detection

### v2

@@ III-B Modal-coordinated Perception Attention (MPA)
The differences in feature distribution between RGB and thermal modalities stem from their distinct physical meanings, encapsulating complementary information. Specifically, while objects share the same spatial location across both modalities, they exhibit varying degrees of saliency, leading to differences in spatial representation. The RGB modality provides richer appearance and texture details, whereas the thermal modality emphasizes overall shape and structure. As a result, the importance of feature channels differs across modalities [19]. This observation highlights the need to align and integrate bi-modal features in both spatial and channel dimensions.
@@
The diagram of MPA is shown in Fig. 2, which follows a Transformer-like structure of Transformer. It primarily consists of a Modal-coordinated Perception Filter (MPF), normalization layers $Norm(\cdot)$, a $3\times 3$ depth-wise separable convolution layer $DConv_{3}(\cdot)$ in the feed-forward network, and residual connections. This process can be formulated as:
@@
where $+$ is element-wise addition. $\widetilde{{f_{i}}}$ and $f_{i}\in\mathbb{R}^{C\times H\times W}$ represent the output of MPF and MPA, respectively, with $C$, $H$, and $W$ denoting the number of channels, height, and width, respectively.
In MPF, we apply a sequence of operations to the normalized bi-modal features $\widetilde{{r_{i}}}$ and $\widetilde{{t_{i}}}$: a $1\times 1$ convolution layer $Conv_{1}(\cdot)$, followed by a StarReLU [75] layer $SR(\cdot)$, and finally, a fast Fourier transform (FFT) in the spatial dimension, to obtain the spatial Fourier components $f_{i}^{rs}$ and $f_{i}^{ts}$:
$f_{i}^{rs}=FFT(SR(Conv_{1}(\widetilde{{r_{i}}}))),$ (3)
$f_{i}^{ts}=FFT(SR(Conv_{1}(\widetilde{{t_{i}}})).$ (4)
In the channel Fourier component enhancement stage, we extract the amplitude $a_{i}^{rs}$ of $f_{i}^{rs}\in\mathbb{R}^{C\times H\times(\frac{W}{2}+1)}$ and $a_{i}^{ts}$ of $f_{i}^{ts}\in\mathbb{R}^{C\times H\times(\frac{W}{2}+1)}$ for the secondary FFT in the channel dimension (CFFT), obtaining the channel Fourier components $f_{i}^{rc}$ and $f_{i}^{tc}$:
$a_{i}^{rs}=Amp(f_{i}^{rs}),a_{i}^{ts}=Amp(f_{i}^{ts})$ (5)
@@
We then enhance the amplitude and phase components using $CLC_{1}(\cdot)$, which consists of two $Conv_{1}(\cdot)$ and a LeakyReLU layer. This process is applied to the amplitude component $a_{i}^{rc}$, the phase component $p_{i}^{rc}$ of $f_{i}^{rc}\in\mathbb{R}^{(\frac{C}{2}+1)\times H\times(\frac{W}{2}+1)}$ and $a_{i}^{tc}$, $p_{i}^{tc}$ of $f_{i}^{tc}\in\mathbb{R}^{(\frac{C}{2}+1)\times H\times(\frac{W}{2}+1)}$, respectively. After that, the channel-enhanced components are composed into complexes and undergo an inverse CFFT (ICFFT). The process is denoted as:
Figure 2: Architectures of the Modal-coordinated Perception Attention (MPA) and its core component Modal-coordinated Perception Filter (MPF).
@@
To effectively integrate the channel-enhanced components with the original amplitude components, we utilize a shortcut connection in each branch. Subsequently, we combine the amplitude and phase components of the spatial Fourier representations from both branches as complexes. We then employ an element-wise addition to facilitate complementarity of two modalities:
$\widetilde{f_{i}^{rs}}=Complex\left({\left({a_{i}^{rs}\cdot\widetilde{a_{i}^{rs}}}\right),CLC_{1}\left({p_{i}^{rs}}\right)}\right),$ (11)
$\widetilde{f_{i}^{ts}}=Complex\left({\left({a_{i}^{ts}\cdot\widetilde{a_{i}^{ts}}}\right),CLC_{1}\left({p_{i}^{ts}}\right)}\right),$ (12)
@@
In the spatial Fourier component enhancement stage, we introduce a dynamic filter $DF\left(\cdot\right)\in\mathbb{R}^{C\times H\times(\frac{W}{2}+1)}$ [66] to perform spatial filtering for the channel-enhanced fusion feature $\widetilde{{f_{i}^{s}}}$. $DF$ consists of $N$ learnable filters $\left\{{{LF_{1}},...,{LF_{N}}}\right\}$ and a multi-layer perception $MLP\left(\cdot\right)$ is used to assign weights for learnable filters. Additionally, we utilize $\widetilde{{r_{i}}}$ and $\widetilde{{t_{i}}}$ to provide initial weights for filters. This process can be formulated as:
$\widetilde{{f_{i}}}=IFFT\left({DF\left({{\widetilde{{r_{i}}}},{\widetilde{{t_{i}}}}}\right)\cdot\widetilde{{f_{i}^{s}}}}\right),$ (14)
@@
where $j$ represents the channel index, $m$ and $n$ denote filter indexes. In this paper, $N$ is set to 8 to avoid over-computing in all experiments.
@@ III-C Frequency-decomposed Edge-aware Block (FEB)
The precise identification of object details is crucial for improving detection performance. However, low-level features extracted from the initial layers of a network often contain boundary details alongside cluttered background noise and interference. This phenomenon poses a significant challenge for the decoder in making accurate predictions.
Figure 4: Architecture of the Edge Frequency Extraction Block (EFEB).
To clarify boundary details and mitigate the impact of noise, we propose the FEB, a novel module specifically designed to provide the decoder with reliable edge features.
In Fig. 3, we present a decomposition of RGB and thermal images in the frequency domain. It can be observed that the amplitude and low-frequency components predominantly capture style and content information, whereas the phase and high-frequency components encapsulate detailed edges and textures. Furthermore, compared to RGB images, thermal images exhibit clearer object contours and less complex textures. Building on this observation, we develop a key component of the FEB, namely the Edge Frequency Extraction Block (EFEB), as illustrated in Fig. 4.
The FEB contains two EFEBs, which begin with $\left\{{{t_{i}},{f_{i}}}\right\}_{i=1}^{2}$ undergoing a Phase Enhancement Process (PEP) to highlight object boundaries, followed by integration. An adaptive high-pass filter $HF$ is then employed to retain detailed edge information while removing style information. Finally, the output is refined by a Fourier Residual Channel Attention Block (FRCAB) to mitigate the neural network’s preference for low-frequency information, with details provided in Sec. III-D.
In PEP, a given input $x$ undergoes a FFT in the spatial dimension initially, followed by phase component enhancement. The PEP can be mathematically expressed as:
@@
$\widetilde{{e_{i}}}=(PEP(t_{i})+PEP(f_{i}))\times HF,$ (19)
${e_{i}}=FRCAB\left({IFFT\left({\widetilde{{e_{i}}}}\right)+{f_{i}}}\right),i=1,2,$ (20)
where $\times$ is element-wise multiplication. Finally, the output edge feature $e_{3}$ of FEB is calculated as:
@@
where $UP_{2}(\cdot)$ represents 2$\times$ up-sampling, and $Concat\left(\cdot\right)$ denotes concatenation. $DNRU(\cdot)$ represents a set of operations including $DConv_{3}(\cdot)$, $Norm(\cdot)$, a ReLU layer, and $UP_{2}(\cdot)$.
@@ III-D Fourier Residual Channel Attention Block (FRCAB)
The FRCAB integrates two key components: the residual-in-residual (RIR) structure [76] and Fourier channel attention (FCA), as shown in Fig. 5. The RIR enables the decoder to focus more on high-frequency information, while FCA extracts inter-channel statistics to further enhance the model’s capability to discriminate global information.
As illustrated in Fig. 5, given an input feature $x\in{{}^{C\times H\times W}}$, the FRCAB operates as follows: First, a global vector is extracted using global average pooling $GAP(\cdot)$ after processing $x$ with a $CLC_{3}(\cdot)$ with $3\times 3$ kernel size. Then, a fast Fourier transform in the channel dimension (CFFT) is performed and $CLC_{1}(\cdot)$ operations are used to enhance the amplitude $a^{x}\in{{}^{(\frac{C}{2}+1)\times 1\times 1}}$ and phase $p^{x}\in{{}^{(\frac{C}{2}+1)\times 1\times 1}}$ components of $x$. Subsequently, an inverse CFFT (ICFFT) is performed, and two residual connections are constructed. Finally, we employ a $DNRU(\cdot)$ to reduce the number of channels and up-sampling features, and obtain the output $y$. The FRCAB can be mathematically described as:
@@
${a^{xc}}={a^{x}}\times GAP\left({CLC_{1}\left(x\right)}\right),$ (24)
${p^{xc}}={p^{x}}\times GAP\left({CLC_{1}\left(x\right)}\right),$ (25)
$x^{ic}=ICFFT\left({Complex\left({{a^{xc}},{p^{xc}}}\right)}\right)\times CLC_{3}\left(x\right)+x,$ (26)
$y=DNRU(x^{ic}).$ (27)
where $\times$ is element-wise multiplication.
Figure 5: Architecture of the Fourier Residual Channel Attention Block (FRCAB).
@@ III-E Bi-domain Learning
We trained our DFENet with various learning strategies in the spatial and frequency domains, aiming to generate accurate saliency maps and reliable edge features.
1) Frequency Domain Learning: As the model deepens, important visual details can be lost, and spectral bias may hinder the model’s ability to learn high-accuracy edge features. To address these problems, we propose Co-focus Frequency Loss (CFL), designed to provide effective guidance for FEB (Sec. III-C). CFL leverages fast Fourier transform (FFT) to measure distances in Fourier domain while cross-referencing details in the original feature space. This allows for the optimization of global information and aids FEB in addressing difficult frequencies.
@@
$\mathcal{E}=UP_{2}\left({DConv_{3}({e_{3}})}\right),$ (34)
and $E$ denotes the edge of objects in ground-truth $G$, obtained by the Canny operator [77].
To focus learning on hard frequencies, we introduce the Co-focus Frequency Matrix (CFM) $w$ to enhance their associated weights. As discussed in Sec. III-C, the phase component contains plentiful edge frequencies. Thus, while adopting $r_{1}$ and $t_{1}$ as co-directors, we set amplitude components of $r_{1}$, $t_{1}$ to the mean values to emphasize their phase components. CFM is expressed as:
@@
~${\widetilde{{\mathcal{R}}}}=Complex(mean\left({Amp\left({FFT\left({{\mathcal{R}}}\right)}\right)}\right),Pha\left({FFT\left({{\mathcal{R}}}\right)}\right))$ (36)
${\widetilde{{\mathcal{T}}}}=Complex(mean\left({Amp\left({FFT\left({{\mathcal{T}}}\right)}\right)}\right),Pha\left({FFT\left({{\mathcal{T}}}\right)}\right))$ (37)
$w=$ (|~$\left(\left|{\widetilde{{\mathcal{R}}}}-FFT(E)\right|+\left|{\widetilde{{\mathcal{T}}}}-FFT(E)\right|\right.$ (38)
$\left.+\left|FFT(\mathcal{E})-FFT(E)\right|\right)^{\alpha},$
@@
2) Spatial Domain Learning: We employ ${{\cal L}_{BCE}}$ to supervise the learning of $\mathcal{E}$ in the spatial domain:
@@
$+{{\cal L}_{IoU}\bigl{(}UP\bigl{(}DConv_{3}(d_{i})\bigr{)},G\bigr{)}}.$
@@ IV Experiments
IV-A Experimental Setup
1) Datasets: To demonstrate the effectiveness of our proposed DFENet, we conducted a series of experiments on four RGB-T SOD benchmark datasets, namely VT821 [11], VT1000 [12], VT5000 [13], and VI-RGBT1500 [14]. The VT821 dataset consists of 821 pairs of RGB-T images captured under different environmental conditions. The VT1000 dataset comprises 1000 aligned RGB-T images with detailed annotations tailored for evaluation, providing a reliable benchmark for performance comparison. The VT5000 dataset offers 5000 pairs of high-resolution, low-bias RGB-T images that cover a broad range of challenging scenes. The VI-RGBT1500 dataset consists of 1500 RGB-T image pairs captured under different illuminations to validate the benefits of incorporating thermal modality. Following previous works [22, 61], we used 2500 RGB-T image pairs from the VT5000 dataset as our training set, while the remaining images and those from the other three datasets served as our testing set.
2) Evaluation Metrics: To comprehensively evaluate the results of each experiment, we employed six widely used evaluation metrics. The E-measure [78] ($Em$) considers both image-level statistics and local pixel differences between the predictions and the ground-truths.

### v3

@@ III-B Modal-coordinated Perception Attention (MPA)
The differences in saliency distribution between RGB and thermal modalities in the frequency domain stem from their distinct physical properties, encapsulating complementary information. Specifically, while objects share the same spatial location across both modalities, they exhibit varying degrees of saliency, leading to differences in spatial representation. The RGB modality provides richer appearance and texture details whereas the thermal modality emphasizes overall shape and structure. As a result, the importance of feature channels differs across modalities. This observation highlights the need to align and integrate bimodal features in both spatial and channel dimensions [21, 36].
@@
The diagram of MPA is shown in Fig. 3, which follows a Transformer-like structure. It primarily consists of a Modal-coordinated Perception Filter (MPF), normalization layers $Norm(\cdot)$, a $3\times 3$ depth-wise separable convolution layer $DConv_{3}(\cdot)$ in the feed-forward network, and residual connections. This process can be formulated as:
@@
where ++ is element-wise addition. $\widetilde{{f_{i}}}$ and $f_{i}\in\mathbb{R}^{C\times H\times W}$ represent the output of MPF and MPA, respectively, with $C$, $H$, and $W$ denoting the number of channels, height, and width.
In MPF, we apply a sequence of operations to the normalized bimodal features $\widetilde{{r_{i}}}$ and $\widetilde{{t_{i}}}$: a $1\times 1$ convolution layer $Conv_{1}(\cdot)$, followed by a StarReLU [83] layer $SR(\cdot)$, and finally, a FFT in the spatial dimension, to obtain the spatial Fourier components $f_{i}^{rs}$ and $f_{i}^{ts}$:
[…]
$).\displaystyle f_{i}^{ts}=FFT(SR(Conv_{1}(\widetilde{{t_{i}}}$)). (4)
To efficiently integrate bimodal features and enhance saliency representation, we decompose the spatial and channel Fourier components of each modality into amplitude and phase, and enhance them individually. As shown, Fig. 4 illustrates the frequency component decomposition of bimodal images where the amplitude captures the intensity distribution and visual saliency of each modality, aiding in the identification of overall salient regions. Enhancing the amplitude component reinforces global saliency cues and suppresses background noise. In contrast, the phase encodes fine-grained spatial details and object boundaries, which are essential for accurate edge delineation and precise localization. Enhancing the phase component refines local structural information and promotes inter-modal structural coherence. This approach enables better preservation of salient cues within each modality and effectively promotes their synergy during fusion.
First, in the channel Fourier component enhancement stage, we extract the amplitude $a_{i}^{rs}$ of $f_{i}^{rs}\in\mathbb{R}^{C\times H\times(\frac{W}{2}+1)}$ and $a_{i}^{ts}$ of $f_{i}^{ts}\in\mathbb{R}^{C\times H\times(\frac{W}{2}+1)}$ for the secondary FFT in the channel dimension (CFFT), obtaining the channel Fourier components $f_{i}^{rc}$ and $f_{i}^{tc}$:
$a_{i}^{rs}=Amp(f_{i}^{rs}),a_{i}^{ts}=Amp(f_{i}^{ts}),$ (5)
@@
We then enhance the amplitude and phase components using $CLC_{1}(\cdot)$, which consists of two $Conv_{1}(\cdot)$ and a LeakyReLU layer. This process is applied to the amplitude component $a_{i}^{rc}$, the phase component $p_{i}^{rc}$ of $f_{i}^{rc}\in\mathbb{R}^{(\frac{C}{2}+1)\times H\times(\frac{W}{2}+1)}$ and $a_{i}^{tc}$, $p_{i}^{tc}$ of $f_{i}^{tc}\in\mathbb{R}^{(\frac{C}{2}+1)\times H\times(\frac{W}{2}+1)}$, respectively.
Figure 3: Architectures of the Modal-coordinated Perception Attention (MPA) and its core component Modal-coordinated Perception Filter (MPF). The RGB feature $r_{i}$ and thermal feature $t_{i}$ serve as inputs to the MPA, producing the output $f_{i}$, while $\widetilde{f_{i}}$ denotes the output of the MPF.
After that, the channel-enhanced amplitude and phase components are composed into complex numbers $Complex(\cdot)$ and undergo an inverse CFFT (ICFFT). The process is denoted as:
@@
To effectively integrate the channel-enhanced components with the original amplitude components, we utilize a shortcut connection in each branch. Subsequently, we combine the amplitude and phase components of the spatial Fourier representations from both branches as complex numbers. We then employ an element-wise addition to facilitate the complementarity of the two modalities:
$\widetilde{f_{i}^{rs}}=Complex\left({\left({a_{i}^{rs}\times\widetilde{a_{i}^{rs}}}\right),CLC_{1}\left({p_{i}^{rs}}\right)}\right),$ (11)
$\widetilde{f_{i}^{ts}}=Complex\left({\left({a_{i}^{ts}\times\widetilde{a_{i}^{ts}}}\right),CLC_{1}\left({p_{i}^{ts}}\right)}\right),$ (12)
@@
where $\times$ is element-wise multiplication.
Second, in the spatial Fourier component enhancement stage, we introduce a dynamic filter $DF\left(\cdot\right)\in\mathbb{R}^{C\times H\times(\frac{W}{2}+1)}$ [75] to perform spatial filtering for the channel-enhanced fusion feature $\widetilde{{f_{i}^{s}}}$. $DF$ consists of $N$ learnable filters $\left\{{{LF_{1}},...,{LF_{N}}}\right\}$ and a multi-layer perception $MLP\left(\cdot\right)$ is used to assign weights for learnable filters. Additionally, we utilize $\widetilde{{r_{i}}}$ and $\widetilde{{t_{i}}}$ to provide initial weights for filters. This process can be formulated as:
[…]
@@
where $j$ represents the channel index, $m$ and $n$ denote filter indexes. In this paper, $N$ is set to 8 to avoid over-computing.
@@ III-C Frequency-decomposed Edge-aware Block (FEB)
The precise identification of object details is crucial for improving detection performance. However, low-level features extracted from the initial layers of a network often contain boundary details alongside cluttered background noise and interference. This phenomenon poses a significant challenge for the decoder in making accurate predictions. To clarify boundary details and mitigate the impact of noise, we propose the FEB, a novel module specifically designed to provide the decoder with reliable edge features.
In Fig. 4, in addition to amplitude reflecting intensity distribution and phase encoding spatial details as discussed in Sec. III-B, we also observe that low-frequency components primarily convey style and content information, while high-frequency components encapsulate detailed edges and textures. Furthermore, RGB images’ phase and high-frequency components contain more complex background textures together with boundary details, whereas those of thermal images exhibit clearer object contours and simpler textures [41, 61, 21, 55]. Building on this observation, we develop a key component of the FEB, namely the Edge Frequency Extraction Block (EFEB), as illustrated in Fig. 5.
Figure 5: Architecture of the Edge Frequency Extraction Block (EFEB). It enhances the low-level thermal feature $t_{i}$ and fusion feature $f_{i}$ through the Phase Enhancement Process (PEP) and an adaptive high-pass filter, thereby obtaining clear edge features $e_{i}$.
The FEB contains two EFEBs. Considering that RGB features contain more background texture noise, EFEBs introduce low-level thermal and fusion features $\left\{{{t_{i}},{f_{i}}}\right\}_{i=1}^{2}$ undergoing a Phase Enhancement Process (PEP) to highlight object boundaries, followed by integration. An adaptive high-pass filter $HF$ is then employed to retain detailed edge information while removing style information. Finally, the output is refined by a Fourier Residual Channel Attention Block (FRCAB) to mitigate the network’s preference for low-frequency information, with details provided in Sec. III-D.
In PEP, a given input $x$ undergoes a FFT in the spatial dimension initially, followed by phase component enhancement. The PEP $PEP(\cdot)$ can be mathematically expressed as:
@@
$\widetilde{{e_{i}}}=IFFT((PEP(t_{i})+PEP(f_{i}))\times HF),$ (19)
${e_{i}}=FRCAB(\widetilde{{e_{i}}}+f_{i}),i=1,2.$ (20)
Finally, the output edge feature $e_{3}$ of FEB is calculated as:
@@
[…]
@@ III-D Fourier Residual Channel Attention Block (FRCAB)
The FRCAB integrates two key components: a Residual Connection (RC) [84] and Fourier Channel Attention (FCA), as shown in Fig. 6. The RC enables the decoder to focus more on high-frequency information, while FCA extracts inter-channel statistics in the Fourier domain to further enhance the model’s capability to discriminate global information.
As illustrated in Fig. 6, given an input feature $x\in\mathbb{R}^{C\times H\times W}$, the FRCAB operates as follows: First, a global vector is extracted using global average pooling $GAP(\cdot)$ after processing $x$ with a $CLC_{3}(\cdot)$ with $3\times 3$ kernel size.
Figure 6: Architecture of the Fourier Residual Channel Attention Block (FRCAB). It mainly consists of a Residual Connection (RC) and a Fourier Channel Attention (FCA), where $x$ and $y$ denote its input and output, respectively.
Then, a CFFT is performed and $CLC_{1}(\cdot)$ operations are used to enhance the amplitude $a^{x}\in\mathbb{R}^{(\frac{C}{2}+1)\times 1\times 1}$ and phase $p^{x}\in\mathbb{R}^{(\frac{C}{2}+1)\times 1\times 1}$ components of $x$. Subsequently, an ICFFT is performed, and two residual connections are constructed. Finally, we employ a $DNRU(\cdot)$ to reduce the number of channels and up-sampling features, and obtain the output $y$. The FRCAB $FRCAB(\cdot)$ can be mathematically described as:
@@
${a^{xc}}={a^{x}}\times CLC_{1}(GAP\left({CLC_{3}\left(x\right)}\right)),$ (24)
${p^{xc}}={p^{x}}\times CLC_{1}(GAP\left({CLC_{3}\left(x\right)}\right)),$ (25)
$\widetilde{{x^{c}}}=ICFFT\left({Complex\left({{a^{xc}},{p^{xc}}}\right)}\right),$ (26)
$y=DNRU(\widetilde{{x^{c}}}\times CLC_{3}(x)+x).$ (27)
@@ III-E Bi-domain Learning
We trained our FreqSal with various learning strategies in the spatial and frequency domains, aiming to generate accurate saliency maps and reliable edge features.
1) Frequency Domain Learning: As the model deepens, important visual details can be lost, and spectral bias may hinder the model’s ability to learn high-accuracy edge features. To address these problems, we propose Co-focus Frequency Loss (CFL), designed to provide effective guidance for FEB (Sec. III-C). CFL leverages FFT to measure distances in the Fourier domain while cross-referencing details in the original feature space. This allows for the optimization of global information and aids FEB in addressing difficult frequencies.
@@
$\mathcal{E}=DConv_{3}({e_{3}}),$ (34)
and $E$ denotes the edge of objects in ground-truth $G$, obtained by the Canny operator [85]. $e_{3}$ is the output of FEB and $\mathcal{E}$ is the predicted edge map.
To focus learning on hard frequencies, we introduce the Co-focus Frequency Matrix (CFM) $w$ to enhance their associated weights. As discussed in Sec. III-C, the phase component contains plentiful edge frequencies. Thus, while adopting $r_{1}$ and $t_{1}$ as co-directors, we set the amplitude components of $r_{1}$, $t_{1}$ to the mean values $Mean(\cdot)$ to emphasize their phase components:
@@
${\widetilde{{\mathcal{R}}}}=Complex(Mean\left({Amp\left({FFT\left({{\mathcal{R}}}\right)}\right)}\right),Pha\left({FFT\left({{\mathcal{R}}}\right)}\right)),$ (36)
${\widetilde{{\mathcal{T}}}}=Complex(Mean\left({Amp\left({FFT\left({{\mathcal{T}}}\right)}\right)}\right),Pha\left({FFT\left({{\mathcal{T}}}\right)}\right)).$ (37)
Then, CFM is expressed as:
$w=$ $(|{\widetilde{{\mathcal{R}}}}-FFT(E)|+|{\widetilde{{\mathcal{T}}}}-FFT(E)|$ (38)
$+|FFT(\mathcal{E})-FFT(E)|)^{\alpha},$
@@
2) Spatial Domain Learning: We also employ ${\cal L}_{BCE}$} to supervise the learning of the edge map $\mathcal{E}$ in the spatial domain:
@@
This bi-domain learning strategy ensures obtaining the high-quality edge map.
@@
$+{{\cal L}_{IoU}\bigl(UP\bigl(DConv_{3}(d_{i})\bigr),G\bigr)},$
where $\left\{{{d_{i}}}\right\}_{i=1}^{4}$ are decoder features generated by FRCABs.
@@ IV Experiments
IV-A Datasets
To evaluate the effectiveness of our proposed FreqSal, we first conducted experiments on the RGB-T SOD task. To further evaluate its generalization, we extended the comparison to RGB-D-T and RGB-D SOD tasks.
1) RGB-T SOD: We conducted experiments on the four RGB-T SOD benchmark datasets, namely VT821 [12], VT1000 [13], VT5000 [14], and VI-RGBT1500 [15]. The VT821 dataset consists of 821 pairs of RGB-T images captured under different environmental conditions. The VT1000 dataset comprises 1000 aligned RGB-T images with detailed annotations tailored for evaluation, providing a reliable benchmark for performance comparison. The VT5000 dataset offers 5000 pairs of high-resolution, low-bias RGB-T images that cover a broad range of challenging scenes. The VI-RGBT1500 dataset consists of 1500 RGB-T image pairs captured under different illuminations to validate the benefits of incorporating thermal modality. Following [15, 41], we used 2500 RGB-T image pairs from the VT5000 dataset as our first training set, while the remaining images and those from the VT821, VT1000, and VI-RGBT1500 datasets served as the testing set.
2) RGB-D-T SOD: We employed the VDT-2048 dataset [86] for the experiments. It provides 2048 aligned RGB-depth-thermal image pairs for robot visual awareness under various light conditions. Following [86], we employed 1048 image pairs from the VDT-2048 dataset as the training set, and the remaining 1000 image pairs formed the testing set.
3) RGB-D SOD: We performed comparison experiments on five RGB-D SOD benchmark datasets, namely NLPR [7], NJUD [8], DUT-RGBD [9], SIP [10], and STERE [11]. The NLPR dataset consists of 1000 RGB-D image pairs containing diverse indoor-outdoor natural scenes. The NJUD dataset contains 1985 image pairs generated from Internet videos and images. The DUT-RGBD dataset contains 1200 RGB-D image pairs from complex real-world scenes. The SIP dataset collects 929 high-definition RGB-D image pairs of human life scenes. The STERE dataset consists of 1000 Internet RGB-D image pairs and corresponding manually labeled masks. Following [57, 27, 24], we conducted two training setups: (1) using 700 image pairs from the NLPR dataset and 1485 from the NJUD dataset, and (2) extending the first setup with an additional 800 image pairs from the DUT-RGBD dataset.
IV-B Evaluation Metrics
To comprehensively evaluate the results of each experiment, we employed six widely used evaluation metrics. The E-measure [87] ($Em$) considers both image-level statistics and local pixel differences between the predictions and the ground-truths. The S-measure [88] ($Sm$) evaluates the structural similarity between the predictions and the ground-truths. The F-measure [89] ($F_{\beta}$}) is a comprehensive metric, calculated by weighted harmonic averaging on precision and recall. We also used the weighted F-measure [90] ($wF_{\beta}$}) to complement the F-measure, addressing issues such as interpolation defects and dependency defects.